Find a cubic function with the given zeros.

-2, 5, -6

f(x) = x3 + 3x2 + 28x - 60

f(x) = x3 - 3x2 - 28x - 60

f(x) = x3 + 3x2 - 28x - 60

f(x) = x3 + 3x2 - 28x + 60


help me please
1
Expert's answer
2014-06-09T03:21:12-0400
If a is a zero of some function then this function divides on (x-a). Of (x-a) is a multiplier of this function. So we get
f(x) = (x+2)(x-5)(x+6) = x3 + 3x - 28x - 60.